### v1

# The Roots of Performance Disparity in Multilingual Language Models: Intrinsic Modeling Difficulty or Design Choices?

Abstract
Multilingual language models (LMs) promise broader NLP access, yet current systems deliver uneven performance across the world’s languages. This survey examines why these gaps persist and whether they reflect intrinsic linguistic difficulty or modeling artifacts. We organize the literature around two questions: do linguistic disparities arise from representation and allocation choices (e.g., tokenization, encoding, data exposure, parameter sharing) rather than inherent complexity; and which design choices mitigate inequities across typologically diverse languages. We review linguistic features, such as orthography, morphology, lexical diversity, syntax, information density, and typological distance, linking each to concrete modeling mechanisms. Gaps often shrink when segmentation, encoding, and data exposure are normalized, suggesting much apparent difficulty stems from current modeling choices. We synthesize these insights into design recommendations for tokenization, sampling, architectures, and evaluation to support more balanced multilingual LMs.
The Roots of Performance Disparity in Multilingual Language Models:
Intrinsic Modeling Difficulty or Design Choices?
Chen Shani1, Yuval Reif2, Nathan Roll1, Dan Jurafsky1, Ekaterina Shutova3 1Stanford University, 2The Hebrew University of Jerusalem, 3University of Amsterdam Correspondence: email@domain
1 Introduction
Multilingual LMs have expanded NLP’s reach by enabling a single model to perform tasks across many languages. They are pretrained on text from hundreds of languages, sharing parameters and representations (Devlin et al., 2019; Conneau et al., 2020; Scao et al., 2022; Imani et al., 2023; Dang et al., 2024). This enables cross-lingual transfer, where patterns learned in one language improve performance in others (Pires et al., 2019; Wu and Dredze, 2019; Lauscher et al., 2020; Malkin et al., 2022a; Blevins et al., 2024). Despite these advantages, persistent performance disparities across languages limit the practical reach of multilingual models Wang et al. (2025); Ghosh et al. (2025); Chang et al. (2024b).
These disparities systematically follow cross-linguistic patterns: higher-resource languages and those structurally similar to dominant training languages generally perform better than low-resource or typologically distant ones Zhao et al. (2025); Akindotuni (2025). The disparities often persist even with large-scale pretraining, suggesting that scaling alone cannot ensure equitable performance (Kaplan et al., 2020; Hoffmann et al., 2022; He et al., 2025). This raises a central question: are some languages inherently harder to model, or do performance gaps reflect engineering artifacts and design choices?
We review how linguistic structure interacts with multilingual design choices to shape performance gaps via two questions: whether disparities stem from intrinsic difficulty or modeling artifacts (e.g., tokenization, data allocation, shared-parameter interference); and which design choices mitigate inequities. We consolidate our findings into a set of recommendations for tokenization, data sampling, model architectures, and evaluation, highlighting where evaluations confound learnability with tokenization or encoding artifacts (Table˜1).
Our synthesis suggests that cross-linguistic gaps rarely reflect intrinsic modeling complexity. Instead, they arise via three mechanisms: (1) shared-parameter training induces negative transfer when typological diversity exceeds effective capacity (Pfeiffer et al., 2022; Blevins et al., 2024; Chang et al., 2024b); (2) tokenization and encoding fragment words or penalize byte-heavy scripts, inflating sequence length without added meaning (Rust et al., 2021; Arnett et al., 2024; Lundin et al., 2025; Land and Arnett, 2025); and (3) data sampling and evaluation misrepresent semantic exposure. Gaps shrink when normalizing segmentation, encoding, and exposure or explicitly allocating capacity, indicating that difficulty stems from modeling choices.
This first systematic review of cross-linguistic modeling difficulty research offers practical design recommendations for multilingual LMs to achieve balanced performance across diverse languages.
Linguistic Factor Observed Artifact Modeling Mechanism Design Levers Orthography and encoding granularity (§2.1) Encoding inefficiency (byte premium); reduced effective exposure under fixed budgets UTF-8 byte-length asymmetries inflate sequence length and reduce effective training signal for many non-Latin scripts Byte-normalized sampling; script-balanced vocabulary construction; character models (§3.2) Morphology: productivity and compounding (§2.2, §2.3) Tokenization (over-segmentation; longer sequences) Frequency-based subwords misalign with morpheme/compound boundaries, yielding inconsistent tokenization and fragmented cues Morphology-aware tokenization; hybrid subword+byte schemes; script- or language-specific vocabularies (§3.1) Information density and redundancy (§2.5) Unequal semantic coverage under fixed token budgets Token-based budgeting allocates unequal information per unit of training, confounding cross-language comparisons Information-, byte- or morpheme-normalized sampling; adaptive scaling of effective exposure to language data (§3.2, §3.4) Typological and syntactic divergence (§2.6, §2.4) Negative transfer under shared parameters; degraded syntactic generalization Shared capacity induces gradient conflict and representation collapse when languages differ strongly in structure Modular capacity and language adapters; typology-aware routing; controlled sharing (§3.4, §3.5) Evaluation sensitivity to tokenization and encoding (§2.1, §2.5) Perplexity comparability confounded by segmentation and byte length Subword-level metrics conflate segmentation decisions with predictability Report character/morpheme-level metrics; add tokenization diagnostics and typology-aware probes (§3.3)
Table 1: Linking linguistic properties to multilingual modeling artifacts, mechanisms, and design levers; section references point to the supporting evidence discussed in the survey.
2 Linguistic Properties
Human languages evolve under multiple, sometimes competing objectives, producing systematic trade-offs across morphology, syntax, and phonology (Gibson et al., 2019). Information may be densely packed within words or distributed across syntax; flexible word order can be balanced by overt marking such as case or agreement. These features preserve overall communicative efficiency despite wide typological diversity (Gibson et al., 2019; Levshina, 2021; Lian et al., 2023). Human acquisition aligns with this view: children reliably acquire their ambient language, though the timing and difficulty of specific constructions vary by typology rather than defining a universal hierarchy of “hard” languages (Slobin, 1987; Berman, 2014).
We review linguistic properties associated with cross-linguistic performance variation, where learnability refers to sample efficiency and predictive performance (perplexity, downstream accuracy). Drawing on NLP, computational linguistics, typology, and information theory, we show how these properties influence tokenization, data allocation, and architecture in multilingual LMs. Each subsection defines a property, summarizes evidence, and discusses factors affecting modeling success.
2.1 Orthography
Orthography drives cross-linguistic disparities by skewing encoding efficiency under equal token budgets.
Orthography describes how languages map linguistic units to written symbols and how these symbols are encoded for modeling. Languages differ in orthographic granularity and transparency, which affects surface predictability even when underlying meanings are comparable (Bright and Daniels, 1996; Katz and Frost, 1992; Ziegler and Goswami, 2005). While these are properties of writing systems rather than inherent linguistic difficulty, orthography directly determines how text is encoded into bytes, characters, or subword units for LMs.
We focus on how orthography shapes modeling through three linked pathways: representational granularity and orthographic depth, encoding length under fixed budgets (bytes per character), and shared-vocabulary tokenization dynamics that favor frequent patterns. Orthographic granularity (e.g., alphabetic, abjad, abugida, syllabary, logographic) affects information density and segmentation stability: logographic scripts pack more information per character, whereas alphabetic scripts distribute it across longer sequences. These pathways interact in multilingual training. Languages that share a script with high-resource corpora benefit from greater subword sharing and more stable segmentation, while scripts with fewer frequent character n-grams yield fewer multi-character subwords and lower information per token (Petrov et al., 2023; Ahia et al., 2023).
Encoding choices also shape shared-vocabulary construction. When BPE operates over UTF-8 bytes, merges start at the byte level rather than characters, disadvantaging scripts with multi-byte characters (Sennrich et al., 2016; Zouhar et al., 2023; Kargaran et al., 2024). Adapting pretrained multilingual models to unseen scripts can introduce widespread UNK tokens, motivating script-aware adaptation (Pfeiffer et al., 2021).
UTF-8 byte-length asymmetry creates a byte premium: Latin characters often require one byte, whereas scripts such as Devanagari, Arabic, or Chinese may require up to three (Yergeau, 2003; Lemire and Muła, 2022; Hilal and Hilal, 2019). In fixed-size corpora, this reduces effective character exposure for non-Latin scripts (Arnett et al., 2024), and for languages like Chinese, Japanese, or Korean, UTF-8 encoding can triple sequence length, shrinking context windows (Moon et al., 2025). Frequency-based subword tokenization further amplifies disparities: high-resource Latin scripts benefit from larger, more informative subwords, while non-Latin scripts are split into shorter fragments, lowering information-per-token (Sennrich et al., 2016; Petrov et al., 2023; Ahia et al., 2023).
Byte- and character-level models reduce subword mismatch but do not fully equalize efficiency: byte premiums still lengthen sequences, and character granularity varies by script (e.g., Chinese characters often align more closely with morphemes than Latin letters), producing different generalization trade-offs (Clark et al., 2022; Xue et al., 2022). Consequently, these approaches trade coverage and fairness for longer sequences and higher compute. Recent methods, such as MYTE and Script-BPE, aim to preserve segment consistency while mitigating script penalties (Limisiewicz et al., 2024; Land and Arnett, 2025).
Orthographic transparency affects surface predictability. Transparent orthographies (Finnish, Turkish) yield predictable sequences, while opaque ones (English, French) increase unpredictability (though vs. through); yet these effects are smaller than those of tokenization (Katz and Frost, 1992; Seymour et al., 2003; Schmalz et al., 2015).
Orthography shapes cross-linguistic disparities through encoding efficiency. Equal token budgets underallocate capacity to non-Latin scripts, though script- or byte-normalized training mitigates this.
2.2 Morphological Complexity
Performance gaps associated with rich morphology stem from segmentation quality, encoding inefficiencies, and data allocation.
Morphological complexity refers to how languages encode grammatical and semantic information within word forms via inflection, derivation, and compounding (Haspelmath and Sims, 2013). While human learners acquire their ambient language reliably (Slobin, 1987; Berman, 2014), morphology has been assumed to increase language modeling difficulty (Cotterell et al., 2018; Gerz et al., 2018; Park et al., 2021; Mielke et al., 2019).
A common explanation is sparsity: morphologically rich languages realize each lemma in many forms, lowering per-form frequency and increasing sample complexity even when morphological rules are regular (Park et al., 2021). For instance, Cotterell et al. (2018) found that decreased LM performance on 21 languages in Europarl correlates with morphological richness, an effect largely removed by lemmatization. Similarly, Gerz et al. (2018) observed substantial morphology effects across 50 typologically-diverse languages.
However, later analyses show that these correlations often reflect modeling artifacts rather than inherent difficulty. Large-scale multilingual studies demonstrate that apparent morphology effects are confounded with tokenization quality, dataset size, and encoding inefficiencies (Mielke et al., 2019). Morphology-aware segmentation (e.g., Morfessor) substantially reduces surprisal gaps induced by standard BPE (Park et al., 2021; Mager, 2022; Saleva and Lignos, 2021), and analyses of WordPiece/BPE segmentations show inadequacies for complex morphology and derivation (Klein and Tsarfaty, 2020; Lerner and Yvon, 2025).
Recent work further shows that controlling tokenization, encoding, and effective exposure largely removes morphology effects. Arnett and Bergen (2025) find that morphological typology (agglutinative, fusional, isolating) does not reliably predict LM performance, whereas byte-length and tokenization disparities can induce spurious correlations (Xue et al., 2022; Rust et al., 2021; Arnett et al., 2024).
Thus, while morphological richness correlates with observed performance gaps, the causal chain appears to run through three mechanisms: 1) segmentation quality determines whether morphemes are preserved or fragmented (Park et al., 2021; Creutz and Lagus, 2007; Saleva and Lignos, 2021; Gutherz et al., 2023); 2) script and byte-encoding inefficiencies increase sequence length for the same semantic content (Bostrom and Durrett, 2020; Arnett et al., 2024; Land and Arnett, 2025; Foroutan et al., 2025); 3) effective data allocation (fixed token budgets and vocabulary learning dynamics) can underexpose rich-morphology and byte-heavy languages (Muller et al., 2021; Chung et al., 2023; Chang et al., 2024b; He et al., 2025).
Consequently, morphology is best treated as an interaction effect: the same typological feature can look harmful under one tokenizer or training budget and largely vanish under another. When these factors are held constant or corrected (e.g., morphology-aware tokenizers, byte-normalized sampling, or language-specific vocabularies), the gap between morphologically simple and complex languages becomes substantially smaller.
2.3 Lexical Diversity and Vocabulary Size
Lexical diversity effects reflect tokenization misalignment, not linguistic complexity.
Lexical diversity captures how many distinct lexical types (lexemes and multiword expressions) a corpus contains and how evenly their frequencies are distributed. In human language acquisition, learning is strongly frequency-driven: low-frequency items are acquired later and remain harder to access, reflecting the long-tailed (Zipfian) distribution of words (Linders and Louwerse, 2022; Zipf, 1935). This connects lexical diversity to learnability: larger effective vocabularies entail longer tails of rare items, raising sample complexity for learning word meanings even if speakers ultimately master them.
Cross-linguistic differences in lexical diversity reflect lexicalization choices (what is expressed as a single word versus a multiword expression) and word-formation productivity (derivation and compounding) (Booij, 2005; Baayen, 2009; Booij, 2010). For instance, languages differ in how motion events are lexicalized (e.g., encoding manner versus path in the verb) (Talmy, 2000; Allen et al., 2007). Lexical diversity is typically measured from word-segmented corpora via type-frequency distributions, using indices like Type-Token Ratio and its length-normalized variants (Covington and McFall, 2010; McCarthy and Jarvis, 2010; Kettunen, 2014; Fergadiotis et al., 2015; Bestgen, 2024). [^1] In corpus linguistics, these indices typically treat “tokens” as word tokens in a word-segmented corpus (not subword tokens produced by NLP tokenizers).
In multilingual LM analyses, lexical diversity predicts perplexity, transfer quality, and output-side measures of generation quality (Mielke et al., 2019; Dehouck and Denis, 2018; Pelloni et al., 2022). However, perplexity alone does not reveal which linguistic attributes are learned (Meister and Cotterell, 2021). Output-side analyses also relate lexical diversity in generations to output quality: Guo et al. (2025) evaluate model outputs along lexical, syntactic, and semantic diversity dimensions and report that higher-quality generations tend to exhibit higher lexical diversity.
More generally, lexical diversity is a robust predictor of difficulty for LMs: Head-POS entropy and raw type counts can outperform typological features (Mielke et al., 2019; Dehouck and Denis, 2018), and tokenization-sensitive measures of the long tail (e.g., Subword Evenness and vocabulary overlap) predict cross-lingual transfer and multilingual perplexity (Pelloni et al., 2022; Limisiewicz et al., 2023). Vocabulary-richness features also predict GPT-2 perplexity and interact with segmentation choices across typologies (Miaschi et al., 2021; Parra, 2024).
However, much of this effect reflects segmentation artifacts: in compounding-prone languages, frequency-based subwords fragment long compounds into many pieces, inflating sequence length and reducing effective exposure per unit of semantic content (Pelloni et al., 2022; Lundin et al., 2025). When segmentation is more consistent or when byte/character models remove subword segmentation, lexical-diversity effects weaken (Arnett et al., 2025). Lexical diversity, therefore, challenges LMs mainly under tokenization schemes that misalign with linguistic structure. Vocabulary size remains a strong predictor, mainly due to segmentation and data sparsity rather than inherent lexical complexity.
2.4 Syntactic Features
Syntactic features affect modeling difficulty indirectly through interactions with morphology, vocabulary size, and tokenization.
Syntactic features describe how languages organize words into phrases and clauses, including word order, case marking, and dependency structure. Syntax and morphology often provide alternative encodings for the same grammatical distinctions: a language may rely more on word order or more on overt marking (case, agreement) to signal roles and relations while preserving overall communicative efficiency (Sinnemäki, 2008; Lian et al., 2023; Levshina, 2021; Fedzechkina et al., 2017; Mollica et al., 2021). In humans, these trade-offs influence which cues must be tracked rather than creating a global difficulty hierarchy.
For LMs, syntactic variation affects surprisal and perplexity, but effects are typically smaller and more indirect than morphology or tokenization (Mielke et al., 2019; Botha and Blunsom, 2014; Limisiewicz et al., 2024; Miaschi et al., 2021; Pimentel et al., 2021). Feature-based analyses typically find that syntactic typology explains less variance than tokenization or lexical measures, with the largest effects occurring when critical syntactic cues rely on morphemes that subword tokenizers fragment (Mielke et al., 2019; Park et al., 2021).
Case marking illustrates this interaction: under standard BPE, languages with productive case systems show higher surprisal, but morphology-aware segmentation reduces the gap by segmenting case morphemes more consistently, increasing their effective frequency and preserving cues for syntactic roles (Park et al., 2021). Word order effects are mixed: basic order alone is not a reliable predictor of perplexity (Mielke et al., 2019), and reducing word-order-specific encoding can improve cross-lingual adaptation (Liu et al., 2021). Probing studies show that models trained on rigid word-order languages rely on positional cues and struggle more with free-word-order languages, where morphological cues must be preserved (Gulordava et al., 2018; Ravfogel et al., 2021). Dependency length adds another constraint: longer dependencies increase context requirements and interact with architectures biased toward English-like branching (Gibson, 1998; Futrell et al., 2015; Hewitt and Manning, 2019).
Syntactic differences rarely cause difficulty in isolation; they interact with tokenization artifacts that inflate sequence length or obscure morphological cues (Arnett and Bergen, 2025). Consequently, syntax-related performance gaps often reflect architectural constraints and English-centric positional heuristics rather than inherent difficulty. Models trained on rigid word order can over-rely on positional shortcuts that transfer poorly to free-word-order or case-rich languages, especially when morphology is fragmented (Gulordava et al., 2018; Ravfogel et al., 2021). Cognitively-motivated inductive biases, such as relative position encodings and syntactically informed attention, can mitigate these issues (Shaw et al., 2018; Dufter et al., 2022; Strubell et al., 2018; Kuribayashi et al., 2024), but positional design choices matter and multilingual evidence for newer schemes (ALiBi, RoPE) remains mixed (Ravishankar and Søgaard, 2021; Press et al., 2022; Su et al., 2024).
Overall, syntactic variation shapes cross-linguistic gaps mainly through interactions with morphology, tokenization, and vocabulary size; once normalized, syntax alone explains less of the variance, though it remains important for generalization and cross-lingual transfer.
2.5 Information-Theoretic Measures
Entropy differences largely capture representational choices, like word length or morphological encoding, rather than the intrinsic learnability of a language.
Information-theoretic metrics quantify predictability and redundancy, but they also reflect morphology, orthography, and other representational choices rather than pure learnability. Some potentially informative metrics remain difficult to define or measure, leaving room for future work. Information-theoretic measures quantify predictability and redundancy: entropy captures average uncertainty, surprisal measures the negative log probability of an observed unit, and compression rate approximates achievable code length under efficient encoding. These metrics provide a principled way to compare languages in terms of predictability and coding efficiency, linking cross-entropy in LMs to fundamental data statistics (Shannon, 1948). In human processing, surprisal theory formalizes the connection between predictability and cognitive difficulty (Hale, 2001; Smith and Levy, 2013).
A central insight from psycholinguistics and quantitative linguistics is that languages maintain stable information rates through compensatory trade-offs. Spoken languages converge on near-constant bits-per-second rates (Coupé et al., 2019; Jaeger, 2010), and morphologically rich languages exhibit higher per-word entropy because they encode more information per word (Bentz et al., 2017; Koplenig and Wolfer, 2023; Koplenig et al., 2025).
Large-scale studies show systematic differences in entropy at the character and word levels, balanced by structural features like word length (Koplenig and Wolfer, 2023; Koplenig et al., 2025). This reflects Uniform Information Density (UID), where languages spread information to keep local surprisal relatively stable (Jaeger, 2010; Jaeger and Levy, 2006), though UID is not a universal law (Meister et al., 2021).
For LMs, entropy interacts with tokenization and sampling: high-entropy sequences require more data, and token-based budgets can exaggerate difficulty when scripts or tokenizers inflate sequence length. Byte-inefficient scripts and fragmented tokenization can inflate apparent entropy without adding semantic content (Rust et al., 2021). At the human-processing level, LM surprisal estimates can predict reading times across multiple languages, suggesting that surprisal is a useful–but imperfect–proxy for cognitive difficulty in cross-linguistic comparisons (Levy, 2008; Goodkind and Bicknell, 2018; Hollenstein et al., 2021; de Varda and Marelli, 2022; Wilcox et al., 2023; Kuperman et al., 2025; Liu et al., 2025).
Controlling for encoding efficiency and tokenization substantially reduces cross-linguistic surprisal gaps and narrows perplexity differences, indicating that part of the observed entropy variation reflects representation and sampling confounds (Arnett et al., 2024; Rust et al., 2021; Foroutan et al., 2025; Tsvetkov and Kipnis, 2024). However, perplexity remains an imperfect proxy for downstream performance: low perplexity can coexist with weak robustness, particularly in low-resource settings (Luitel et al., 2025; Gurgurov et al., 2025; Zhuang and Sun, 2025; Liu et al., 2022; Lourie et al., 2025).
Compression-based metrics provide architecture-independent baselines by evaluating predictability at fixed representational units. Bits-per-character/byte (BPC) estimates cross-entropy per character/byte, reducing reliance on subword tokenization and enabling comparisons that align with LM perplexity and transfer (De Souza et al., 2024; Tsvetkov and Kipnis, 2024). However, BPC is encoding-sensitive: UTF-8 byte premiums and script granularity distort comparisons even after byte normalization (Arnett et al., 2024; Moon et al., 2025; Foroutan et al., 2025; Deletang et al., 2024).
In sum, information-theoretic differences reflect language encoding choices rather than inherent learnability, and normalizing for density, byte length, or morphemes reduces many cross-linguistic gaps.
2.6 Typological Distance
Typological diversity causes difficulty only in shared-parameter & -vocabulary settings.
Moving from single-language difficulty to cross-linguistic transfer, typological distance captures similarity in grammar (syntax, morphology), lexicon (cognates, word choice), and phonology, whereas genealogical relatedness reflects shared ancestry. Linguistic diversity represents alternative solutions to similar communicative constraints, with no single feature reliably predicting processing difficulty (Bickel, 2013; Comrie, 1989; Ponti et al., 2019). In human L2 acquisition, linguistic distance predicts attainment and learning difficulty, though effects are often mediated by specific domains such as phonology, morphology, or lexicon (Chiswick and Miller, 2005; Isphording and Otten, 2014; Schepens et al., 2020).
Typological similarity and relatedness also shape cross-lingual transfer: shared ancestry and structural resemblance determine when parameter sharing helps or hurts. Early multilingual models show that shared vocabularies bias representations toward related languages, with mBERT organizing languages along genealogical lines (Pires et al., 2019; Wu and Dredze, 2019; Rama et al., 2020). Transfer is strongest among closely related languages with overlapping lexicon and morphology. For example, Dutch models transfer better to German than to English despite English’s larger data volume (de Vries et al., 2021; Muller et al., 2023). Relatedness can also increase interference: near-identical forms can induce negative transfer (Lauscher et al., 2020).
At a finer level, WALS-based similarity predicts transfer quality beyond raw resource size (Dryer and Haspelmath, 2013; Lin et al., 2019; Malkin et al., 2022a), with features like word order and head direction particularly predictive (K et al., 2020; Blaschke et al., 2025). Tokenization-based diagnostics (e.g., Subword Evenness, vocabulary overlap) and information-theoretic metrics (Information Parity) can predict cross-lingual transfer (Pelloni et al., 2022; Tsvetkov and Kipnis, 2024; Limisiewicz et al., 2023).
These effects are pronounced in large-scale multilingual LMs. The curse of multilinguality refers to declining per-language performance as more languages share parameters (Conneau et al., 2020; Gurgurov et al., 2024), with low-resource and typologically distant languages suffering most (Lauscher et al., 2020). Typological distance amplifies interference and exacerbates vocabulary fragmentation. Controlled studies show that adding related languages improves low-resource performance, but distant languages harm both settings as shared capacity becomes overburdened (Chang et al., 2024b). Gradient conflicts are common when distant languages are trained jointly (Wang et al., 2020).
Overall, typological and morphological analyses converge on a single insight: shared-parameter training can induce interference as typological diversity grows, and transfer may weaken when shared vocabularies fragment the long tail in morphologically rich languages (Arnett and Bergen, 2025; Park et al., 2021; Chang et al., 2024b), but Kallini et al. (2025) found that every overlap helps. Modular approaches that allocate language-specific capacity reduce conflict while preserving positive transfer (Pfeiffer et al., 2022; Blevins et al., 2024).
2.7 Summarizing Linguistic Properties
Across features, many performance gaps arise from mismatches between linguistic structure and modeling choices rather than intrinsic language difficulty. Tokenization and encoding can fragment cues and lengthen sequences, sampling can create unequal exposure, and shared-parameter training can cause negative transfer when typological diversity exceeds capacity. These factors also confound evaluation: low perplexity does not guarantee robust downstream performance, especially in low-resource settings. When segmentation, encoding, and exposure are normalized, many apparent cross-linguistic gaps shrink, showing that current modeling paradigms, not linguistic diversity itself, drive much of the disparity.
3 Design Implications
These findings motivate design implications for tokenization, sampling, architecture, evaluation, and corpus construction; we focus on interventions most directly supported by the surveyed evidence.
3.1 Tokenization: From Frequency-Based Segments to Linguistically Informed Units
Tokenization is a widely studied design lever. Frequency-driven subword algorithms (BPE, WordPiece) often fragment morphology and non-Latin scripts, inflating sequence length, training costs, and obscuring linguistically meaningful units (Park et al., 2021; Mielke et al., 2019; Ali et al., 2024a; Lundin et al., 2025; Gutherz et al., 2023; Ali et al., 2024b). Across studies, segmentation quality explains a substantial portion of cross-linguistic gaps, where morphology- and script-aware methods reduce surprisal in agglutinative and polysynthetic languages (Creutz and Lagus, 2007; Ling et al., 2015; Mager, 2022; Saleva and Lignos, 2021; Loukatou et al., 2021; Dang et al., 2025).
Operationalizing tokenization quality remains nontrivial: widely used diagnostics include compression/sequence length and corpus token count, alongside distributional measures such as Rényi-entropy-style vocabulary balance, with mixed evidence on when compression improvements translate to model quality (Gallé, 2019; Schmidt et al., 2024; Gutherz et al., 2023; Chizhov et al., 2024; Goldman et al., 2024; Dagan et al., 2024). Poor tokenization harms low-resource settings (Nag et al., 2025), but engineering interventions can offset training overheads (Hong et al., 2024).
Implication 1: Use byte-/character-level models, and treat script-specific tokenizers as first-class components (Qarah and Alsanoosy, 2024; Alrefaie et al., 2024; Rana et al., 2025). Adaptive and multilingual approaches that balance scripts and allocate shared-vocabulary capacity can reduce parity gaps without compromising high-resource performance (Bostrom and Durrett, 2020; Limisiewicz et al., 2023; Ahia et al., 2024; Baxi and Bhatt, 2024). Byte-level and tokenizer-free models show competitive performance at scale, trading longer sequences for reduced script penalties (Kenter et al., 2018; Pagnoni et al., 2025). We note that byte-level models are more expensive, and thus might be less usable for a fixed budget.
3.2 Data Sampling and Byte Normalization
Token-based sampling penalizes byte-heavy scripts, whereas byte-normalized sampling narrows gaps (Arnett et al., 2024; Wei et al., 2021).
Parity metrics highlight large variability in tokenization quality across scripts (Kanjirangat et al., 2025), motivating sampling strategies that target semantic exposure rather than raw token counts (e.g., UniMax, byte-premium scaling;  Chung et al., 2023; Chang et al., 2024a; He et al., 2025).
Implication 2: Pretraining should use byte-normalized, information-normalized, or morpheme-normalized sampling for equal semantic coverage across languages. Data balancing should reflect linguistic diversity rather than corpus availability, correcting for segmentation bias, type proliferation, and script inefficiency.
3.3 Beyond One-Size-Fits-All Benchmarks
Current multilingual benchmarks often conflate linguistic difficulty with tokenization artifacts or dataset size. Perplexity is sensitive to tokenizer choice, whereas character-, morpheme-, and byte-level metrics provide more robust comparisons (Xue et al., 2022; Clark et al., 2022; Tsvetkov and Kipnis, 2024; Kanjirangat et al., 2025; Fang et al., 2024). Tokenizer-quality diagnostics and standardized reporting help disentangle measurement bias from true modeling capability (Chelombitko et al., 2024; Bender and Friedman, 2018).
Cross-linguistic syntactic challenge suites like CLAMS offer controlled tests of generalization and reveal consistent gaps between monolingual and multilingual models (Mueller et al., 2020). For morphology, community benchmarks (e.g. SIGMORPHON) provide fine-grained metrics that complement perplexity-based ones (Cotterell et al., 2017).
Probe-based evaluations show representational disparities, such as weaker subject/object identification in case-rich languages when models are trained on fixed word order (Ravfogel et al., 2021; Papadimitriou et al., 2021), motivating typology-aware competency assessments.
Implication 3: Evaluation should use linguistically informed metrics and typology-aware probes beyond subword perplexity. Benchmarks should disaggregate performance by morphology, script, and word order to avoid masking inequities.
3.4 Balanced Corpora and Pretraining
Pretraining corpus choice strongly shapes multilingual performance. Web data overindexes English and underrepresents high-vitality languages, correlating poorly with global populations (Dunn, 2020; Dunn and Adams, 2020; Mehmood et al., 2017; Mor, 2025; Joshi et al., 2020; Khanna and Li, 2025; Bella et al., 2023). Corpus composition often tracks speaker counts over linguistic diversity, while data statements support accountable multilingual reporting (Bender and Friedman, 2018).
Multilingual corpora favor Indo-European languages and underrepresent complex morphology, minority scripts, and small populations. Balancing must account not only for token counts but also for linguistic density: information per token, morphological productivity, and rare-form distributions.
Resources such as UniMorph and high-coverage dependency treebanks can support typology-aware evaluation of coverage, even without direct training supervision (Nivre et al., 2020). These findings motivate moving beyond a single monolithic model: leveraging language similarity or tailoring components to typologically related clusters can boost learning for low-resource languages without forcing uniform representations (Malkin et al., 2022b).
Implication 4: Corpus design should explicitly encode linguistic diversity by accounting for representational efficiency and linguistic density, ensuring that languages with high morphological or typological variation receive equivalent semantic coverage, not merely equivalent token counts.
3.5 Linguistic “Difficulty” as a Constraint
Ultimately, “difficulty” in multilingual LMs reflects mismatches between model assumptions and linguistic structure. Cross-linguistic disparities shrink when units are preserved, sampling normalized, and architectures interference-free (Jaeger, 2010; Gibson et al., 2019; Bickel, 2013).
Implication 5: Treat linguistic diversity as a design constraint, ensuring model components respect language structure. This framing supports typology-aware models that respect linguistic diversity rather than reinforcing English-centric biases.
4 Conclusions and Future Work
Multilingual performance is shaped less by inherent linguistic complexity than by design choices: tokenization, data allocation, and interference-aware training. Future work should explore language-adaptive strategies: predicting data and capacity needs per language, designing curricula that prioritize transfer from related languages, and developing architectures that dynamically allocate resources across typologically distinct languages. Truly low-resource and endangered languages require innovative approaches under scarcity.
Evaluation must also evolve: metrics should reflect cross-linguistic differences in task difficulty while capturing fairness and accessibility. Aligning model inductive biases with human learning can guide more robust multilingual NLP.
By embracing linguistic diversity as a design principle, we can build models that are more adaptable, equitable, and capable of supporting the full spectrum of the world’s languages.
5 Limitations
Despite our analysis of multilingual performance, several limitations warrant consideration. First, our work focuses primarily on representation- and architecture-driven factors (tokenization, encoding, shared parameters) and does not fully capture other potential sources of difficulty, such as pragmatic, discourse-level, or sociolinguistic phenomena, which may affect real-world usage.
Second, most of our empirical insights rely on pretrained models and standard evaluation datasets, which may underrepresent truly low-resource or endangered languages. Data sparsity, orthographic variation, and non-standardized corpora in such languages could yield patterns not observed in higher-resource languages.
Third, while we consider cross-linguistic typology, our analysis is largely English-centric in architecture and benchmark design, which may bias conclusions about syntax, word order, and positional encoding effects.
Fourth, information-theoretic measures capture correlations with morphology and orthography rather than intrinsic learnability. Metrics for hierarchical structure, discourse-level predictability, or multimodal signals remain underexplored, leaving important aspects of language modeling outside our current framework.
Finally, despite our thorough literature survey, it is possible that relevant works were overlooked. We welcome pointers to such papers to keep this survey up to date.
Addressing these limitations in future work will be crucial for building truly language-adaptive, equitable, and robust multilingual models.
AI usage: The paper used AI assistance for rephrasing, for finding additional relevant papers, and occasionally for summarizing them.